A 21.5 mL sample of 0.383 M dimethylamine, (CH32NH, is titrated with 0.279 M perchloric acid. (1) Before the addition of any perchloric acid, the pH is (2) After adding 11.8 mL of perchloric acid, the pH is (3) At the titration midpoint, the pH is (4) At the equivalence point, the pH is (5) After adding 42.2 mL of perchloric acid, the pH is

Ka of dimethylamine = 5.9*10^-4
Answer #1

(a)

(CH3)2NH is a weak base.

Ka of (CH3)2NH = 5.9 * 10^-4

So, Kb = Kw / Ka
            = (1.0 * 10^-14) / (5.9 * 10^-4)
            = 1.7 * 10^-11

Now,

                 (CH3)2NH    +    H2O      <------->        (CH3)2NH2+     +     OH-
IC:               0.383                                                        0                      0
C:                 - x                                                          + x                    + x
EC:          0.383 – x                                                      x                      x

Kb = [(CH3)2NH2+] [OH-] / [(CH3)2NH]
or, 1.7 * 10^-11 = (x) (x) / (0.383 – x)
or, 1.7 * 10^-11 = x2 / (0.383 – x)
or, 1.7 * 10^-11 = x2 / (0.383)          [Kb is very small, so x term in denominator is discarded]
or, x2 = (1.7 * 10^-11) * (0.383)
or, x2 = 6.5 * 10^-12
or, x = 2.55 * 10^-6

So, [OH-] = x = 2.55 * 10^-6

pOH = - log [OH-]
        = - log (2.55 * 10^-6)
        = 5.59

Now, pH = 14 – pOH
              = 14 – 5.59
              = 8.41

(2)

Moles of (CH3)2NH = 0.383 M x 0.0215 L = 0.0082345 mol
Moles of perchloric acid = 0.279 M x 0.0118 L = 0.0032922 mol

So, 0.0032922 mol of perchloric acid will react with 0.0032922 mol of (CH3)2NH to produce 0.0032922 mol of (CH3)2NH2+.

So, mole of (CH3)2NH left = 0.0082345 mol - 0.0032922 mol = 0.0049423 mol

Total volume = 21.5 mL + 11.8 mL = 33.3 mL = 0.033 L

So, the concentration terms are
[(CH3)2NH] = 0.0049423 mol / 0.033 L = 0.1498 M
[(CH3)2NH2+] = 0.0032922 mol / 0.033 L = 0.0998 M

Using the Henderson - Hasselbalch equation

pOH = pKb + log ([salt] / [base])
pOH = pKb + log { [(CH3)2NH2+] / [(CH3)2NH] }

Some preliminary calculations :
Kb = 1.7 * 10^-11

So, = pKb = - log Kb = - log (1.7*10^-11) = 10.77

Now,

pOH = pKb + log { [CH3)2NH2+] / [CH3)2NH] }
pOH = 10.77 + log (0.0998 / 0.1498)
pOH = 10.77 + log (0.666)
pOH = 10.77 - 0.18
pOH = 10.59

Now, pH = 14 – pOH
              = 14 – 10.59
              = 3.41

(3)

At mid point;
Half of the [(CH3)2NH is neutralized by the acid. Therefore, the moles of (CH3)2NH = moles of (CH3)2NH2+.

So, [(CH3)2NH] = [(CH3)2NH2+]

Using the Henderson - Hasselbalch equation

pOH = pKb + log { [(CH3)2NH2+] / [(CH3)2NH] }
pOH = pKb + log 1
pOH = 10.77 +
pOH = 10.77

Now, pH = 14 – pOH
              = 14 – 10.77
              = 3.23

4.

At equivalence point;

At equivalence point, all the base [(CH3)2NH] is neutralized with acid.

Moles of (CH3)2NH = 0.0082345 mol

So, moles of perchloric acid = 0.0082345 mol
molarity of perchloric acid = 0.279 M
So, volume of perchloric acid = 0.0082345 mol / 0.279 M = 0.0295 L = 29.5 mL

Total volume = 21.5 mL + 29.5 mL = 51 mL = 0.051 L

Now, 0.0082345 mol of perchloric acid will react with 0.0082345 mol of (CH3)2NH to produce 0.0082345 mol of (CH3)2NH2+.

[(CH3)2NH2+] = (0.0082345 mol) / (0.051 L) = 0.16 M

Now,

                 (CH3)2NH2+          <------->         (CH3)2NH       +       H+
IC:               0.16                                               0                      0
C:                 - x                                               + x                    + x
EC:          0.16 – x                                              x                      x

Ka = [(CH3)2NH] [H+] / [(CH3)2NH2+]
or, 5.9*10^-4 = (x) (x) / (0.16 – x)
or, 5.9*10^-4 = x2 / (0.16 – x)
or, 5.9*10^-4 = x2 / (0.16)          [Ka is very small, so x term in denominator is discarded]
or, x2 = (5.9*10^-4) * (0.16)
or, x2 = 9.44 * 10^-5
or, x = 0.0097

So, [H+] = x = 0.0097

pH = - log [H+]
        = - log (0.0097)
        = 2.01

(5)

Moles of (CH3)2NH = 0.383 M x 0.0215 L = 0.0082345 mol
Moles of perchloric acid = 0.279 M x 0.0422 L = 0.0117738 mol

So, 0.0082345 mol of CH3)2NH will react with 0.0082345 mol of perchloric acid.

So, mole of perchloric acid left = 0.0117738 mol - 0.0082345 mol = 0.0035393 mol

Total volume = 21.5 mL + 42.2 mL = 63.7 mL = 0.0637 L

So, [perchloric acid] = 0.0035393 mol / 0.0637 L = 0.056 M

Perchloric acid is a strong acid and hence will dissociate completely.

So, [H+] = 0.056 M

pH = - log [H+]
    = - log (0.056)
      = 1.25